Netflix is selling for $105 a share. A Netflix call option with one month until expiration and an exercise price of $116 sells for $2.30 while a put with the same strike and expiration sells for $13.10.

a. What is the market price of a zero-coupon bond with face value $116 and 1-month maturity? (Round your answer to 2 decimal places.)

b. What is the risk-free interest rate expressed as an effective annual yield? (Round your answer to 2 decimal place.)
Answer #1

Using Put Call Parity,

C + X/(1 + r)t = Stock Price + P

2.30 + 116/(1 + r)1/12 = 105 + 13.10

r = 2.09%

a.

Market Price of Bond = 116/(1 + 0.0209)1/12

Market Price of Bond = $115.80

b.

EAR = (1 + 0.0209/12)12 - 1

EAR = 2.11%